OpenStudy (anonymous):

solve for x, where x is a real number: x4−1/x3=0

OpenStudy (anonymous):

\[x^4-1/x^3 = 0\]

OpenStudy (anonymous):

Solve for x over the real numbers: -1/x^3+x^4 = 0 Write the left hand side as a single fraction. Bring x^4-1/x^3 together using the common denominator x^3: (-1+x^7)/x^3 = 0 Multiply both sides by a polynomial to clear fractions. Multiply both sides by x^3: -1+x^7 = 0 Isolate terms with x to the left hand side. Add 1 to both sides: x^7 = 1 Elminate the power on the left hand side. Take 7^th roots of both sides: Answer: | | x = 1